OpenStudy (anonymous):

Which quadratic function does not have the domain and the range y > 6? y = 6(x – 4)(x – 2) y = 0.5(x – 3)2 – 6 y = 0.5x2 – 2x – 4 y = x2 – 2x – 6 would it be the last one ?

OpenStudy (mertsj):

No

OpenStudy (anonymous):

then which one ?

OpenStudy (anonymous):

???/

OpenStudy (mertsj):

And the domain of all of them is all real numbers so I think you have posted something incorrectly.

OpenStudy (anonymous):

yes just one sec i just noticed Which quadratic function does not have the domain X\[\epsilon\] Real numbers and the range y > 6? y = 6(x – 4)(x – 2) y = 0.5(x – 3)2 – 6 y = 0.5x2 – 2x – 4 y = x2 – 2x – 6

OpenStudy (anonymous):

the X equalls to all real number

OpenStudy (anonymous):

ok thank you

OpenStudy (mertsj):

yw. You posted something wrong.

OpenStudy (mertsj):

Actually the last one has range y >=-7